OpenStudy (anonymous):

Which of the following is the solution to the equation 8(y – 2) = 2y ?

OpenStudy (anonymous):

Sorrry its. 8^(y – 2) = 2^y ?

OpenStudy (jhannybean):

\[\large 8^{(y-2)}=2^y\] first we need to make the left hand side have the same base. \[\large \quad\quad 8 = 2^3\]\[\large 2^{3(y-2)}=2^y\] Now that they have the same base,you could eliminate the bases and just solve for the powers. \[\large 3(y-2)=y\]Can you take it from here? :)

OpenStudy (anonymous):

wait so im supposed to solve 3(y-2) ?

OpenStudy (anonymous):

help please someone :?

jimthompson5910 (jim_thompson5910):

yes you need to solve 3(y-2)=y for y

OpenStudy (jhannybean):

\[\large \color{blue}3(y)-\color{blue}3(2)= y\] does that help?

OpenStudy (anonymous):

so 3y-6=y?

OpenStudy (jhannybean):

yes and now you want to bring the y's to one side and your constants on the other.

OpenStudy (anonymous):

2y=6 y=3?

OpenStudy (jhannybean):

subtract the "y" from both sides and add a "6" to both sides.\[\large 3y -6 \color{purple}{+6} \color{green}{-y} = 0\color{red}{+6}\] now simplify both sides. \[\large 3y - y = 6\] can you solve for y now?